### v1

# Multi-Modal Artificial Intelligence of Embryo Grading and Pregnancy Prediction in Assisted Reproductive Technology: A Review

Abstract
As a global disease, infertility has always affected human beings. The development of assisted reproductive technology can effectively solve this disease. However, the traditional in vitro fertilization-embryo transfer technology still faces many challenges in improving the success rate of pregnancy, such as the subjectivity of embryo grading and the inefficiency of integrating multi-modal data. Therefore, the introduction of artificial intelligence-based technologies is particularly crucial. This article reviews the application progress of multi-modal artificial intelligence in embryo grading and pregnancy prediction based on different data modalities (including static images, time-lapse videos and structured table data) from a new perspective, and discusses the main challenges in current research, such as the complexity of multi-modal information fusion and data scarcity.
Index Terms:
artificial intelligence, assisted reproductive technology, multi-modal.List of Abbreviations
| ACC | Accuracy |
| AFC | Antral Follicle Count |
| AI | Artificial Intelligence |
| AMH | Anti-Mullerian Hormone |
| ART | Assisted Reproductive Technology |
| AUC | Area Under receiver operating characteristic Curve |
| BC | BlastCoele |
| BG | Background |
| BMI | Body Mass Index |
| CNN | Convolutional Neural Network |
| EMT | EndoMetrial Thickness |
| GNN | Graph Neural Network |
| Grad-CAM | Gradient-weighted Class Activation Mapping |
| hCG | Human Chorionic Gonadotropin |
| ICM | Inner Cell Mass |
| ICSI | IntraCytoplasmic Sperm Injection |
| IVF-ET | In Vitro Fertilization-Embryo Transfer |
| PGT | Preimplantation Genetic Testing |
| RNN | Recurrent Neural Network |
| SHAP | SHapley Additive exPlanations |
| TE | TrophEctoderm |
| TLM | Time-Lapse Microscopy |
| t-SNE | t-distributed Stochastic Neighbor Embedding |
| ZP | Zona Pellucida |
I Introduction
As women age, especially over the age of 35, their fertility declines and the risk of pregnancy increases. Studies have shown that up to 10-15 of couples of childbearing age are diagnosed with infertility worldwide [1]. It is of great significance for the family and the society to help infertile couples obtain their own children through modern medical technology. For families, long-term infertility often makes both husband and wife have great psychological pressure. The emergence of Assisted Reproductive Technology (ART) provides new hope for them. This can not only greatly improve their psychological state, relieve their fertility pressure, but also enhance the happiness of the family. For the society, the popularization of ART can provide more reproductive choices for families in their countries and regions. Thus, it helps them escape from the fertility dilemma, promotes the reasonable growth of the population, and improves the fertility rate to a certain extent.
In Vitro Fertilization-Embryo Transfer (IVF-ET), commonly known as test-tube baby, is a key technology in the field of assisted reproduction [2]. In this technique, oocyte and sperm are fertilized outside the body to form an embryo, which is then transferred back to the woman’s uterus to achieve pregnancy. According to clinical statistics, the pregnancy success rate of a single IVF-ET transfer cycle is only 40-50, and the final live birth rate is as low as 30 [3]. There are many reasons for this result. First of all, when selecting high-quality embryos to be transferred to the female uterus in clinical practice, it mainly relies on the embryologist’s observation of the embryo image to grade the embryo quality, and the high grade is preferred to transfer. However, this method is highly subjective and error-prone, because the results of multiple assessments by the same embryologist or the same assessment by different embryologists are quite different [4]. In addition, embryo grading has a weak correlation with final pregnancy prediction, so even high-quality embryos selected by experienced embryologists may not result in pregnancy after transfer. Secondly, the key factors affecting pregnancy include not only embryo quality, but also parental fertility indicators, such as parental age, number of high-quality embryos, endometrial thickness, sperm quality and other clinical indicators [5]. Reproductive doctors need to predict the success of pregnancy according to embryo quality and fertility indicators of the couple, which once again introduces subjectivity and uncertainty. Moreover, reproductive doctors and embryologists are not the same group of people, so the coordination between different departments once again affects the outcome of pregnancy. Therefore, the success rate of IVF-ET under the traditional clinical decision-making mode is low, and patients may need to undergo multiple embryo transfers to achieve successful pregnancy, which will lead to an increase in treatment time and cost. In addition, patients who have experienced a failure of IVF-ET treatment are often unwilling to continue to receive more treatment, because the process of ART is complex and the time span is large, which causes great psychological pressure to patients [6].
The above-mentioned problems highlight the urgent need for automated Artificial Intelligence (AI) embryo grading and pregnancy prediction in practical applications. Firstly, the automated system has higher standardization in assessing the quality of embryos, which effectively eliminates the human subjective bias, thus significantly improving the accuracy and reducing the cost of embryo selection. Secondly, automated technologies enable efficient processing of large amounts of embryo data, greatly reducing the workload of embryologists and reproductive doctors, enabling them to focus on handling complex cases and in-depth communication with patients. Finally, AI can not only provide decision support, but also have a certain degree of interpretable analysis ability, which can reveal the key influencing factors behind model predictions, and AI can integrate multi-modal information of patients to achieve highly individualized treatment recommendations. The main advantages of AI in ART are shown in Fig. 1 (a).
II Related work
Dalal et al. [7] systematically introduced the tasks of gamete assessment for sperm and oocytes, and summarized a variety of methods for automated analysis, which laid a foundation for the objective and standardized assessment of gamete quality. Furthermore, Deshpande et al. [8] and Merican et al. [9] focused on the morphological characteristics of embryos and summarized automated embryo grading methods, to improving the reproducibility and predictive efficiency of traditional subjective grading. In particular, Isa et al. focused on the automated analysis of blastocyst stage embryos and summarized a series of AI-based evaluation methods, including key tasks such as blastocyst quality grading and blastocyst image segmentation. From a more macro perspective, Hanassab et al. [10] comprehensively sorted out the research status and development trend of gamete assessment and embryo assessment. Given that the ultimate goal of embryo grading is to optimize transfer success, some researchers are beginning to incorporate pregnancy outcome prediction into the evaluation framework. For example, Mapstone et al. [11] and Louis et al. [12] summarized models for embryo morphology grading and pregnancy outcome prediction.
In addition to the review from the perspective of specific tasks, Dimitriadis et al. [13] proposed a systematic classification method based on the embryonic development stages (including sperm stage, oocyte stage, pronuclear stage, cleavage stage and blastocyst stage), which integrated the automated tasks involved in each stage from the perspective of time. Alsaad et al. [14] focused on time-lapse imaging data and summarized relevant AI algorithms based on its unique temporal information. At the application level, Mapari et al. [15] and Alikani et al. [16] more emphasized the potential of artificial intelligence technology in the clinical scenario of assisted reproduction, especially the possibility of combining robotics and automated operating systems to realize the whole process. In order to further improve the comprehensive understanding of the application of ART, some efforts [17][18][19][20] have attempted to integrate gamete assessment (including sperm and oocytes), embryo assessment (cleavage and blastocyst stage), ploidy prediction, preimplantation genetic testing (PGT), pregnancy prediction and other tasks into the same review framework.
The main tasks of the related work are summarized as shown in Fig. 1 (b). However, these works often divide tasks roughly into several categories, and do not deeply explore the specific effects of different data modalities on the suitability of task modeling and algorithms. In fact, there are many types of data involved in assisted reproductive practice, including static images, time-lapse videos, structured tabulated data, etc. Therefore, reorganizing tasks based on data sources can not only depict the problem more accurately, but also help clarify the rationality and limitations of the model design. This review work is precisely based on this motivation. Focusing on the two main tasks: embryo grading task and pregnancy prediction task in ART, we conduct a systematic sorting and comparison starting from different data modalities. The main contributions of this paper include:
We conducted a systematic review of previous studies on embryo grading and pregnancy prediction, classified by data modalities, and compared the datasets used, modeling methods, and model performance.
We summarized the ART related public datasets and discussed the core challenges in current research, especially the key issues such as multi-modal information fusion and insufficient data in the ART.
Our work not only provides a comprehensive overview of current research on automated embryo grading and pregnancy prediction techniques, but also suggests the way for future research.
III Background
With the development of society and the change of lifestyle, infertility has attracted more and more attention. The emergence and development of ART has provided hope for many couples who are unable to conceive naturally. The birth of the world’s first test-tube baby in 1978 marked the beginning of ART [21]. In recent years, with the continuous progress of ART technology, remarkable achievements have been made in improving the pregnancy rate, reducing complications and personalized treatment, which largely provides more reproductive choices for patients with infertility. In the development process of ART technology, it can be divided into the first generation of IVF-ET, the second generation of IntraCytoplasmic Sperm Injection (ICSI), and the third generation of Preimplantation Genetic Testing (PGT). This section provides a brief introduction to the technology at each stage of ART.
III-A The first generation——IVF-ET
IVF-ET is currently the most widely used method in ART, which mainly targets infertility caused by female factors such as tubal obstruction and ovulation disorder. Louise Brown, the world’s first test-tube baby born through IVF-ET, was born in Britain in 1978, marking a major breakthrough in the field of human assisted reproduction [22]. The process of IVF-ET is shown in the Fig. 2(a).
Step 1: By controlling ovarian ovulation with gonadotropin injection, we can obtain multiple mature oocytes.
Step 2: By collecting sperms and purifying them, we then allow the sperms to combine with oocytes in vitro to obtain zygotes.
Step 3: The zygotes (embryos) are continuously cultured and dynamically monitored in a constant temperature incubator at 37°C, usually for 3 to 5 days.
Step 4: Based on the morphological characteristics, we can select high-quality embryos with high developmental potential.
Step 5: The selected embryo is transferred into the mother’s uterus.
Step 6: Pregnancy is confirmed by three methods. Biochemical pregnancy, about 10 days after embryo transfer, confirms whether the embryo implants by detecting the level of Human chorionic gonadotropin (hCG), which is an early sign of pregnancy. Fetal heart clinical pregnancy, 5 to 6 weeks after embryo transfer, an ultrasound examination is conducted to confirm whether the embryo in the uterus has a fetal heart beating, which is a mid-pregnancy sign. Live birth, whether the newborn shows vital signs after delivery, is the ultimate sign of pregnancy.
In the overall IVF-ET process, high-resolution microscopic imaging can be used to capture continuous embryo development image data in the step 3, and clinical tabular data related to parental fertility can be simultaneously collected from step 1 to step 5. These tabular data include, but are not limited to: Female age, Body Mass Index (BMI), Endometrial Thickness (EMT), Antral Follicle Count (AFC), quality of sperm, etc.
In the step 3, it is very important to continuously observe the image of embryo development for judging embryo quality and improving the success rate [23]. The process of the first 5 days of embryonic development is shown in the Fig. 2(d).
Day 0: The sperm enters the oocyte and the genetic material of the sperm and oocyte combine to form the zygote.
Day 1: The zygote starts the first cleavage, forming a 2-cell embryo.
Day 2: The embryo continues cleavage at the 2-4 cell stage.
Day 3: The embryo continue cleavage at the 6-8 cell stage.
Day 4: The embryo continues to divide, forming a 16-32 cell morula, where cells begin to join tightly and individual cell boundaries become blurred to form a dense cell mass.
Day 5: The morula continues to develop to form the blastocyst. The blastocyst has a distinct BlastoCoel (BC), and the internal cells begin to differentiate to form two main parts: the Inner Cell Mass (ICM), which will develop into the fetus; The TrophEctoderm (TE) develops into the placenta. The Zona Pellucida (ZP) of the blastocyst begins to thin and is ready for hatching and implantation.
In clinical practice, cleavage embryos (day 3) or blastocyst embryos (day 5) are usually used for transfer. Cleavage embryos are used because some patients have a low number of embryos and are concerned about losing embryos to culture at the blastocyst stage. In addition, due to limited culture conditions in some laboratories, the success rate of culture of cleavage embryos is higher than that of blastocyst. Blastocyst embryos are selected because after an additional 2-3 days of culture, embryo quality is naturally screened and inferior embryos are eliminated. In addition, if PGT genetic screening is needed, blastocyst embryos are more suitable because there are more cells in the embryo to sample.
During the IVF-ET process, there are two key human decision-making points, step 4 and step 5 in the Fig. 2(b) and (c), which give rise to the two tasks of embryo grading and pregnancy prediction. For the task of embryo grading, embryologists need to select the embryo with the most developmental potential as the subsequent transfer object according to the characteristics of embryo development, division rate, cell symmetry and so on. For the task of pregnancy prediction, reproductive doctors combine the embryo grading results with parental fertility indicators to comprehensively determine the embryo transfer scheme and optimize the pregnancy success rate. Therefore, embryo grading and pregnancy prediction have become two core tasks in the clinical practice of assisted reproduction and are highly dependent on professional experience. The application of automated decision-making systems in embryo grading and pregnancy prediction has important clinical value and development potential.
III-B The second generation——ICSI
ICSI was first successfully applied in clinical practice in 1992 by a research team from the Free University of Brussels in Belgium. This breakthrough marked an important step in the field of ART. ICSI is mainly used for the treatment of infertility caused by male factors, especially severe oligospermia, asthenospermia, and azoospermia. Although the overall procedure of ICSI is similar to that of conventional IVF-ET, there are significant differences in sperm acquisition and fertilization stages [24]. In the sperm acquisition stage, ICSI has relatively low requirements for sperm quality, and sperm can be obtained by testicular or epididymal puncture even for azoospermia patients. During the fertilization stage, ICSI uses a micromanipulative technique to inject a single sperm directly into the cytoplasm of the egg cell, thereby bypassing the natural sperm penetration through the zona pellucidum and cell membrane of the egg cell. This technique has significantly improved the fertilization success rate of male infertility patients and brought hope to many families.
III-C The third generation——PGT
PGT is a revolutionary technology in the field of ART. PGT uses genetic analysis of early embryos cultured in vitro to screen embryos without specific genetic abnormalities for transfer, aiming to reduce the abortion rate, improve the success rate of pregnancy, and effectively block the transmission of single-gene genetic diseases or chromosomal abnormalities to offspring [25]. PGT is generally performed at the blastocyst stage, usually on day 5-6 after fertilization, and 5 to 10 TE cells are taken from the blastocyst for testing, without involving the ICM that will develop into the fetus in the future. Blastocyst stage biopsy can not only ensure the embryo development potential, but also provide more adequate genetic material for analysis. Through PGT technology, precise genetic screening and disease prevention can be achieved before embryo implantation, which greatly improves the overall success rate of ART.
IV AI in ART
With the continuous development of medical imaging and machine learning technology, AI is gradually applied to the IVF-ET process, bringing new trends to reproductive medicine. In the process of IVF-ET, embryo grading and pregnancy prediction are two key steps that affect the pregnancy outcome. Traditionally, embryo grading mainly relies on manual assessment by experienced embryologists based on the morphological characteristics of embryos, while pregnancy prediction is usually based on comprehensive analysis of parental clinical characteristics (such as age, BMI, quality of sperm) and high-quality embryo image features by reproductive doctors.
As a technology that can automatically learns feature representation and performs predictive analysis from large-scale and complex data, AI has shown important potential in the field of IVF-ET. Through Convolutional Neural Network (CNN), Recurrent Neural Network (RNN) and other methods, researchers can directly extract features from static embryo images, temporal embryo images, and multi-modal clinical information to effectively improve the consistency of embryo grading and the accuracy of pregnancy prediction. It should be pointed out that, as shown in the Fig. 3, the selected methods are not the same for different tasks and different data modality. Based on this, this section systematically reviews the research of AI-based embryo grading and pregnancy prediction in recent years, focusing on the current technical path, model performance, data resources, main challenges and future development trends in clinical application, so as to provide theoretical support and practical reference for the further optimization of intelligent ART.
IV-A Embryo grading
Embryo grading plays a crucial role in ART. By carefully evaluating the morphological characteristics of embryos, this process helps clinicians to select the most promising embryos for transfer, thereby significantly improving the success rate of IVF. Traditionally, embryo grading mainly relies on the manual observation of embryo images by embryologists and a comprehensive score based on a series of morphological indicators, such as cell number, cell size uniformity, integrity of the zona pellucida, and the degree of fragmentation. With the rapid development of computer vision technology, AI-based embryo grading methods have gradually become a research hotspot.
Depending on the type of input data, the embryo grading tasks can be divided into two main categories: grading based on static embryo images and grading based on temporal embryo images. The former is usually evaluated using a single image of an embryo developing to a specific stage, such as day 5. While the latter analyzes image sequences of the embryo at multiple consecutive time points, such as day 1 to day 5, to capture its dynamic development process. As shown in Fig. 3, in the ”Task” column: Figure (a) corresponds to the embryo grading task based on the static cleavage stage image (day 3). Embryos are divided into four grades: A (high), B, C, and D (low). The higher the grade, the more regular the arrangement of cells in the embryo. Figure (b) corresponds to the embryo grading task based on the static blastocyst stage image (day 5). Embryos are classified into two grades: ”Good” (high) and ”Bad” (low), and the higher the grade, the more complete the development of the blastocyst. In addition, the dataset for embryo grading task based on temporal embryo images is shown in Fig. 2(d). The input of the dataset is a series of images, and the labels are classified according to the embryo development level, just like static embryo images.
It is important to note that embryo grading tasks are often closely linked to other related tasks, such as cell counting and embryo segmentation. Embryo development grade is closely related to the number of cells. Therefore, accurate cell counting is essential for embryo grading. At the same time, the embryo segmentation task aims to divide the embryo into different regions, such as the ICM and TE, whose morphological characteristics are of great reference value for assessing the embryo. By analyzing these derived tasks in depth, a more complete understanding of the developmental state of the embryo can be achieved, thereby improving the accuracy and reliability of embryo grading. As shown in Fig. 3, in the ”Task” column: Figure (c) corresponds to the cell counting task; Figure (d) corresponds to the blastocyst segmentation task.
| Paper | Dataset (number, type) | Year | Method | Task | ACC |
| Septiandri[26] | 1084, day3 image (static) | 2020 | ResNet50 | 3-embryo grading | 91.79% |
| Mohamed[27] | 110, day5 image (static) | 2023 | VGG16 | 4-embryo grading | 90.00% |
| Wu[28] | 3601, day3 image (static) | 2021 | CNN+Ensemble | 4-embryo grading | 74.14% |
| Wu[29] | 3599, day3 image (static) | 2022 | CNN+Attention | 4-embryo grading | 82.98% |
| Perumal[30] | 840, day3 and day5 image (static) | 2024 | GAT | 2-embryo grading | 98.80% |
| Liu[31] | 276801, day2-day6 image (static) | 2025 | Transformer | 16-embryo grading | 73.10% |
| You[32] | 5736, day3 image (static) | 2025 | UNet+GCN | 4-embryo grading | 86.89% |
| Guo[33] | 7728, day3 image (static) | 2022 | Grad-CAM+ResNet | 4-embryo grading | 78.61% |
| Guo[34] | 7728, day3 image (static) | 2023 | UNet+CNN | 4-embryo grading | 79.56% |
| Wang[35] | 1025, day5 image (11 focal) | 2021 | VGG16 | 2-embryo grading | 89.90% |
| Liu[36] | 4997, day5 image (7 focal) | 2024 | ResNet50 | 10-embryo grading | 86.53% |
| Thirumalaraju[37] | 1188, day5 image (static) | 2020 | Xception | 5-embryo grading | 64.95% |
| Paya[38] | 3844, day4 and day5 image (static) | 2022 | ResNet50 | 2-embryo grading | 82.77% |
| Abbasi[39] | 127, day5 image (temporal) | 2021 | ResNet18+Concat | 2-embryo grading | 73.08% |
| Abbasi[40] | 130, day3 and day5 image (temporal) | 2021 | CNN+Add | 2-embryo grading | 76.90% |
| Kragh[41] | 4032, day5 image (temporal) | 2019 | CNN+RNN | 6-embryo grading | 67.40% |
| Vaidya[23] | 803, day 2,3 and 5 image (temporal) | 2021 | VGG16+LSTM | 5-embryo grading | 100% |
| Kalyani[42] | 704, day1-day3 image (temporal) | 2024 | ResNet50+GRU | 2-embryo grading | 93.05% |
| Wang[43] | 531, day1-day3 image (temporal) | 2023 | R(2+1)D | 2-embryo grading | 77.74% |
| Fordham[44] | 136, day1-day5 image (temporal) | 2022 | Transformer | 3-embryo grading | 62.50% |
| Malmsten[45] | 661062, day1-day5 image (static) | 2020 | Inception V3 | 8-cell counting | 93.90% |
| Nguyen[46] | 148918, day1-day3 image (static) | 2022 | Transformer | 4-cell counting | 94.30% |
| Zhou[47] | 442, day3 image (static) | 2022 | ResNet50+Genetic algorithm | 20-embryo grading | 86.41% |
IV-A1 Static embryo grading
When grading embryos based on static embryo images, the current processing strategy is generally using CNN to extract image features and then grade embryos. Septiandri et al. [26] used ResNet50 [48] to perform a 3-grading task on the 3rd day of embryo development, and the final classification Accuracy (ACC) could reach 91.79. Similarly, Mohamed et al. [27] used VGG16 [49] to perform a 4-grading task on the 5th day of embryonic development, and the final ACC was 90.00. In order to combine the advantages of different CNNs, Wu et al. [28] collected 3601 images on the 3rd day of embryo development of 1800 couples for 4-grading, and extracted the features of DenseNet169 [50], Inception V3 [51], ResNet50 and VGG19 last layer for ensemble classification, respectively. And the final ACC was 74.14. Attention mechanism, as a module to enhance feature extraction, is often used in combination with CNN to improve the model’s attention to key regions during embryonic development. Wu et al. [29] added channel-level attention and point-level attention to CNN, enabling the model to adaptively focus on regions related to the embryo grade. Eventually, the ACC of embryo 4-grading on the test set was 79.15, and the ablation experiments also demonstrated the effectiveness of the attention mechanism.
In addition to CNN, some recent studies have also begun to introduce Graph Neural Network (GNN) and Transformer [52] into embryo grading tasks. For example, Perumal et al. [30] directly converted images into graph nodes and input them into GAT [53], and extracted features for embryo 2-grading through the message passing mechanism of GNN. Liu et al. [31] converted images into patches, obtained local features through self-attention operation within patches, and obtained global features through Transformer among patches, which were finally used for embryo 16-grading.
The two-stage model is a modeling strategy often used to handle complex tasks. Its core idea is to divide the overall task into two interconnected sub-stages to achieve more efficient and accurate prediction. In the first stage, the model usually performs the initial screening task, aiming to extract potential key regions from the input; In the second stage, the model conducts fine classification on the candidate results generated in the first stage to further optimize the accuracy rate. You et al. [32] and Guo et al. [33][34] extracted the blastocyst segmentation map and Grad-CAM [54] attention map of embryos in the first stage, respectively, so that the model in the second stage could focus on regions of interest with embryo grading.
In addition, since the microscope can take multi-angle embryo images with different focal lengths, some studies have begun to explore methods for embryo grading after fusion of these multi-angle images [35][36]. This multi-angle image fusion strategy is expected to further improve the accuracy and reliability of embryo grading.
IV-A2 Temporal embryo grading
In addition to basic microscopy, Time-Lapse Microscopy (TLM) is often used to observe embryos. This microscope is capable of taking TLM images, which are a set of temporally sequenced embryo images as shown in Fig. 2(d). TLM can capture the changing process of cells over time, showing the process of cell division, morphological change, movement, development, etc. Compared with the static image, TLM images have an additional time dimension. In order to solve the problem of processing multiple time-series images, Abbasi et al. [39][40] extracted the features of a single image by CNN, and directly fused the features of multiple images by addition or concatenation operation for the subsequent embryo grading task. Apart from these simple and direct methods, Kragh et al. [41], Vaidya et al. [23] and Kalyani et al. [42] extracted single image features through Inception V3, VGG16 or ResNet50, and then integrated the features of TLM images through RNN for embryo internal quality grading.
For the embryo grading task, whether it develops into a blastocyst is an important indicator of embryo quality. Wang et al. [43] directly converted the embryo grading task into a blastocyst classification task, and used a multi-task learning method to train two tasks at the same time: multi-frame blastocyst classification task were assisted by the single-frame classification task. On the multi-frame task model, ResNet was used to extract single frame information and then R(2+1)D network [55] (a 3DCNN method) was used to integrate multi-frame information, and the ACC reached 77.74. Due to the natural advantages of Transformer model in processing temporal data, Fordham et al. [44] directly converted temporal images into tokens and input them into Transformer for blastocyst classification, and finally ACC is 62.50.
The above methods based on static or temporal images realize the automation of embryo quality grading, significantly improve the efficiency and accuracy of embryo evaluation, and the grading effect has surpassed that of professional embryologists, which provides a standardized auxiliary tool for embryologists. The details of these methods are shown in the TABLE I.
IV-A3 Cell counting
Although the embryo grading task can effectively evaluate the quality of the embryo, it requires professional embryologists to label the images. However, the label acquisition of embryo cell counting task is easier, and the number of embryo cells can indirectly represent the quality of the embryo, so there are also a considerable number of studies on embryo cell counting. Malmsten et al. [45] used the improved Inception V3 model to count 8 embryonic cells, and the ACC was as high as 93.90. For the same cell counting task, Nguyen et al. [46] not only considered the spatiality of a single image, but also considered the temporality of multiple continuous embryo development images. By simultaneously inputting a sequence of embryo development images into Transformer, the time information was introduced to help the model carry out more accurate embryo cell 4-counting. When performing embryo grading, exfoliated cells around the embryo can interfere with the judgment of automated grading. Therefore, Zhou et al. [47] extracted the region of interest to eliminate exfoliated cells near the embryo during preprocessing, and used adaptive genetic algorithm to find the best parameters of the network, which completed the two tasks of cell counting and embryo grading simultaneously.
| Paper | Dataset | Year | Method | Region of | Dice |
| (number, type) | segmentation | ||||
| Saeedi[56] | 221, day5 image (static) | 2017 | K-means+ | TE | 77.30% |
| Watershed algorithm | ICM | 83.10% | |||
| Harun[57] | 2019 | UNet | TE | 92.50% | |
| ICM | 94.30% | ||||
| Rad[58] | 2020 | Dilated U-Net | TE | 86.61% | |
| Corpuz[59] | 2023 | ResUNet | TE | 91.20% | |
| ICM | 98.70% | ||||
| Chen[60] | 2025 | Embryo-Net | TE | 81.20% | |
| ICM | 86.51% | ||||
| Miled[61] | 2025 | ResA-Net | TE | 81.80% | |
| ICM | 83.09% | ||||
| ZP | 86.87% | ||||
| Rad[62] | 2019 | Blast-Net | TE | 76.52% | |
| ICM | 81.07% | ||||
| BC | 80.79% | ||||
| ZP | 81.15% | ||||
| BG | 94.74% | ||||
| Mushtaq[63] | 2022 | ECS-Net | TE | 78.43% | |
| ICM | 85.26% | ||||
| BC | 88.41% | ||||
| ZP | 85.34% | ||||
| BG | 94.87% | ||||
| Arsalan[64] | 2022 | SSS-Net | TE | 78.15% | |
| ICM | 84.50% | ||||
| BC | 88.68% | ||||
| ZP | 84.51% | ||||
| BG | 95.82% | ||||
| Arsalan[65] | 2022 | MASS-Net | TE | 79.08% | |
| ICM | 85.88% | ||||
| BC | 89.28% | ||||
| ZP | 84.69% | ||||
| BG | 96.07% | ||||
| Ishaq[66] | 2023 | FSBS-Net | TE | 80.17% | |
| ICM | 85.55% | ||||
| BC | 89.15% | ||||
| ZP | 85.80% | ||||
| BG | 95.62% | ||||
| Arsalan[67] | 2024 | PSF-Net | TE | 80.00% | |
| ICM | 86.46% | ||||
| BC | 90.15% | ||||
| ZP | 85.77% | ||||
| BG | 96.10% | ||||
| Jamal[68] | 327, day3 image (static) | 2023 | DenseUNet | Cell | 96.62% |
| Zhang[69] | 1548, day3 image (static) | 2024 | SAM | Fragments | 69.50% |
IV-A4 Embryo segmentation
Image segmentation of embryos at the cleavage stage and blastocyst stage aims to accurately identify and quantify different structures and cell compositions during embryonic development, thereby providing a reliable basis for embryo quality assessment in ART. At the blastocyst stage, segmentation mainly focuses on Inner Cell Mass (ICM), TrophEctoderm (TE), BlastCoele (BC) and Zona Pellucida (ZP) to support more detailed blastocyst scoring and PGT biopsy. Saeedi et al. [56] was the early work that proposed automated embryo segmentation, using K-means and watershed algorithm to segment ICM and TE, and the final Dice scores were 83.10 and 77.30, respectively. In addition, this work published relevant dataset, a total of 211 embryo images with segmentation labels, which promoted the development of this field. Based on the previous work, many studies [57][58][59] began to segment ICM and TE by improving UNet [70]. In addition, Chen et al. [60] proposed a segmentation model based on spatial modeling based on the prior knowledge that the ICM is in the center and the TE is in the periphery. In addition to the segmentation of ICM and TE, there are many studies [61][62][63][64][65][66][67] which have included BC, ZP and background (BG) regions into the segmentation criteria.
Although the segmentation of blastocyst stage embryos is the mainstream, there are also some works segmenting cleavage stage embryos. At the cleavage stage, segmentation helps to extract key morphological features such as cell number, symmetry, and cytoplasmic distribution, which are closely related to the developmental potential of the embryo. Jamal et al. [68] used DenseUNet to segment the cells by collecting and labeling images of 327 embryos on the third day of development, and the final Dice score was 96.62. Zhang et al. [69] introduced a SAM-based [71] two-branch segmentation method, in which the semantic branch was used to segment fragments, and the instance branch was used to detect and segment cleavage stage embryos. The final mAP on the embryo cells was 87.40, and the Dice score on the fragments was 69.50.
The embryonic cell counting task and the embryo segmentation task are two key computer vision tasks in the field of ART. The embryonic cell counting task mainly focuses on the images of embryos at the cleavage stage. The embryo segmentation task focuses on embryo images at the blastocyst stage. Although the two are at different stages of embryonic development and have different analytical goals, they are both committed to achieving the automation and standardization of embryo evaluation. The details of these methods are shown in the TABLE I and TABLE II.
| Paper | Dataset (number, type) | Year | Method | Score |
| Xu[72] | 185, day3 image (static) | 2014 | LBP+SVM | ACC=64.71% |
| Geller[73] | 361, day5 image (static) | 2021 | Inception V1 | AUC=65.70% |
| Sawada[74] | 470, day1-day5 image (static) | 2021 | CNN+Attention | AUC=64.20% |
| Berntsen[75] | 14644, day1-day5 image (temporal) | 2021 | I3D+LSTM | AUC=67.00% |
| Abbasi[76] | 130,day3 and day5 image (temporal) | 2023 | CNN+Transformer | ACC=77.50% |
| Borna[77] | 252, day 1,2,3 image (temporal) | 2024 | UNet+CNN | ACC=75.00% |
| Ouyang[78] | 4046, day1,2,3 image (temporal) | 2025 | Transformer | AUC=61.70% |
| Mapstone[79] | 700, day1-5 image (temporal) | 2024 | MobileNet | AUC=68.00% |
| Liu[80] | 401, 19 tabular data | 2021 | RF | AUC=61.30% |
| Mehta[81] | 333, 11 tabular data | 2024 | AdaBoost | ACC=97.50% |
| Dehghan[82] | 812, 25 tabular data | 2024 | AdaBoost+Genetic algorithm | AUC=91.00% |
| Mohammad[83] | 10036, 46 tabular data | 2025 | RIMARC | AUC=92.12% |
| Sergeev[84] | 8732, 19 tabular data | 2024 | DNN | AUC=79.80% |
| Gong[85] | 657, 25 tabular data | 2023 | LR | AUC=87.90% |
| Zhu[86] | 969, 23 tabular data | 2024 | LR | AUC=76.50% |
| Wu[87] | 15779, 11 tabular data | 2024 | RF | AUC=71.60% |
| Cao[88] | 2865, 123 tabular data | 2024 | GPT | AUC=87.00% |
| Ouyang[78] | 4046, 22 tabular data | 2025 | Transformer | AUC=71.30% |
| Enatsu[89] | 19342, day5 image (static)+13 tabular data | 2022 | ResNet18+RF | AUC=71.00% |
| Charnpinyo[90] | 1099, day5 image (static)+4 tabular data | 2023 | CNN+MLP | AUC=72.00% |
| Kim[91] | 2555, day5 image (static)+1 tabular data | 2024 | Image segmentation+CNN+MLP | AUC=74.10% |
| Liu[92] | 17580, day5 image (static)+16 tabular data | 2022 | CNN+MLP | AUC=77.00% |
| Kim[93] | 3286, day1-day5 image (temporal)+ unknown tabular data | 2024 | Transformer | AUC=68.80% |
| Ouyang[78] | 4046, day1,2,3 image (temporal)+22 tabular data | 2025 | Transformer+DeFusion | AUC=74.60% |
IV-B Pregnancy prediction
Although deep learning has made great progress in the task of embryo grading and has surpassed that of humans, the purpose of embryo grading is for pregnancy prediction. The role of embryo grading in pregnancy prediction is indirect, so current research tends to directly predict pregnancy outcomes. Pregnancy prediction is to predict the success of pregnancy after embryo transfer by analyzing the parents’ clinical data, embryo characteristics data, and laboratory test data. Current AI-based pregnancy prediction tasks can be divided into image-based, table-based and multi-modal based tasks according to different data types. Image-based task can be further divided into static image-based and temporal image-based. One point to be stated is that in this section we classify all three types of pregnancy, hCG biochemical pregnancy, fetal heart clinical pregnancy, and live birth, as pregnancies. As shown in Fig. 3, in the ”Task” column, Figure (e) represents the negative and positive embryos for pregnancy prediction based on the third day or the fifth day.
IV-B1 Pregnancy prediction based on static image
The earlier machine learning method for pregnancy prediction based on static embryo images can be traced back to 2014. Xu et al. [72] extracted the features of static embryo images to 256 dimensions by Local Binary Pattern (LBP) [94]. Then Support Vector Machine (SVM) [95] was used to classify the extracted features, and the comprehensive performance of the test data in six institutions was ACC=64.71. In 2021, Geller et al. [73] collected 361 blastocyst stage embryo images and used the Inception V1 network to predict pregnancy. The final Area Under receiver operating characteristic Curve (AUC) was 65.70. Sawada et al. [74] fused the static embryo image information extracted by CNN branch and attention mechanism branch for pregnancy prediction, and the AUC reached 64.20.
IV-B2 Pregnancy prediction based on temporal image
Previous image-based pregnancy prediction studies generally use images of the last day of embryo development to predict pregnancy, which ignores the timing of embryo development. Therefore, pregnancy prediction studies based on temporal embryo images have been gradually developed. Berntsen et al. [75] used I3D network [96] (a 3DCNN method) to obtain the features of 128 sequences of embryo development images, and then reduced the feature dimension by pooling operation. Finally, bidirectional LSTM [97] was used to connect temporal features for the final pregnancy prediction.
Unlike image analysis, video analysis requires spatial-temporal data analysis, but many methods for processing spatial-temporal data such as 3DCNN are not suitable for medical applications due to high computational complexity. Therefore, Abbasi et al. [76] divided the pregnancy prediction task of temporal embryo image into three stages. In the first stage, CNN was used to extract individual image features on day 3 and day 5, and in the second stage, the temporal image features were fused through the attention mechanism in Transformer. In the third stage, the results of day 3 and day 5 time-series images were integrated by voting method to make the final pregnancy prediction. When processing TLM images, many frames of images will be recorded for each embryo, but these images have a lot of redundant information because the changes of embryos are small in a short time. Borna et al. [77] only used the three key development images of the first, second and third days of embryo development for pregnancy prediction. The ACC of the model reached 75.00, far higher than 60.32 of professional embryologists. Also using keyframe temporal images, Ouyang et al. [78] proposed a spatial-temporal position encoding, making the Transformer more suitable for fusing temporal images.
In order to understand the process of embryo development more comprehensively, Mapstone et al. [79] carried out the embryo grading task and pregnancy prediction task at the same time, and finally the ACC of the embryo 5-grading task was 88.00, and the AUC of pregnancy prediction was 68.00.
IV-B3 Pregnancy prediction based on tabular data
In clinical practice, reproductive doctors need to predict pregnancy outcomes not only based on embryo images, but also based on tabular data such as parental fertility indicators. Tabular data is usually structured and easily quantified, and does not require the collection of large amounts of data like image processing tasks. Liu et al. [80] collected 401 cases of data with 19 fertility indicators, such as female age, female BMI, and EndoMetrial Thickness (EMT), and used SVM, Logistic Regression (LR) [98], Decision Tree (DT) [99], and Random Forest (RF) [100] to predict pregnancy. The AUC of the best RF method was 61.30, and it was found through statistics that: Among patients with good-quality embryos, the clinical pregnancy success rate was about 70 when EMT thickness was greater than 9.6 mm, but 50 when EMT thickness was no greater than 9.6 mm. Mehta et al. [81] collected data of 333 cases with 11 indicators such as female age, Antral Follicle Count (AFC), and Follicle-Stimulating Hormone (FSH), and observed that the ACC of pregnancy prediction of Adaboost [101] was 97.50, better than all other machine learning models, and found statistically that: The propensity to clinical pregnancy is negative if the woman is older than 36 years, and the likelihood of clinical pregnancy decreases substantially if the woman is older than 40 years; Furthermore, the propensity for clinical pregnancy was positively correlated with the number of embryos transferred in the same IVF cycle.
Similarly, through statistical related studies [82][83][84][85][86][87], it can be concluded that the characteristics affecting pregnancy prediction mainly include female age, Anti-Mullerian Hormone (AMH), EMT, semen volume, oocytes number, AFC, female BMI and so on. In addition to using traditional machine learning methods for pregnancy prediction, Cao et al. [88] converted 123 relevant data into natural language and fine-tuned the large language model GPT-4 with the converted natural language, resulting in a final AUC of 87.00. Ouyang et al. [78] converted the tabular data into tokens and input them into Transformer for pregnancy prediction with AUC of 71.30.
IV-B4 Pregnancy prediction based on multi-modal data
With the progress of AI and the improvement of medical data collection technology, pregnancy prediction is more likely to use multi-modal data such as images and tables. This trend is mainly due to the potential of multi-modal data fusion, which integrates the advantages of different types of data to provide more comprehensive prediction results, thereby improving the clinical application value of pregnancy prediction. Compared with image data, tabular data can directly express high-level semantics information. Therefore, some works [89][90] extract image information into low-dimensional data through CNN, and then input it into the fusion model combined with tabular data directly. The quality of the embryo at the blastocyst stage is mainly reflected from two aspects, ICM and TE. Based on this prior knowledge, Kim et al. [91] first binarized and segmented the ICM and TE of the embryo through the traditional image segmentation method, and then added the segmentation map to the original image to obtain the embryo enhancement image. Finally, the features extracted from the enhancement image were combined with age features in the pregnancy prediction task. Similarly, Liu et al. [92] collected 17580 multi-modal data for pregnancy prediction, in which the image data contained two images focusing on ICM and TE respectively, and the tabular data contained 16 parental fertility indicators.
Considering that the image modality data used in the previous multi-modal pregnancy study are all static, Kim et al. [93] collected the time series embryos image data from day 1 to day 5 and health spreadsheet data. When processing the image data, it was multi-stage. Firstly, each morphology of the embryo was segmented through the segmentation model. Then the segmentation images were input into the neural network to predict five key indicators: blastocyst size, blastocyst level, cell boundary, cell number and developmental stage. After processing, the original time series image, spreadsheet data, morphological segmentation image and five key indicators were input into Transformer for final pregnancy prediction. In addition, Ouyang et al. [78] are the first to integrate images of the first 3 days of embryo development and fertility table data for pregnancy prediction, and propose an effective decoupling fusion network to fuse multi-modal data.
In terms of the performance of pregnancy prediction task, models based on tabular data generally significantly outperform models using only images. This may be due to the closer temporal distance and stronger correlation between tabular data and pregnancy outcomes, thus enabling the model to capture more direct and critical predictive features. In addition, tabular data are often extracted by clinical experts with highly correlated features of pregnancy outcomes, which has strong prior value and further improves the performance of the model. It is worth noting that even in the multi-modal modeling scenario, the tabular model still generally outperforms the multi-modal methods that fuse images and tables. This phenomenon may not result from the ineffectiveness of multi-modal fusion itself, but more reflects the realistic challenges existing in current multi-modal research. For example, existing work is mostly based on private datasets, and there are significant differences among different institutions in data collection processes, feature encoding methods, and image device resolutions, which leads to the lack of comparability of model performance under different data sources. The details of these methods are shown in the TABLE III.
| Dataset | Year | Modality | Description | Download link |
| A time-lapse embryo dataset for morphokinetic parameter prediction.[102] | 2022 | Static image | The dataset consists of 704 time-lapse videos of developing embryos with 7 focal length, covering day 2-6 of embryo development, for a total of 2.4 million images. Annotations are graded to 16 different developmental stages. | https://zenodo.org/records/6390798 |
| Embryologist agreement when assessing blastocyst implantation probability: is data-driven prediction the solution to embryo assessment subjectivity?[44] | 2022 | Temporal image | The dataset consists of 143 time-lapse videos of embryo development, including day 1-5 of embryo development. Annotations are 36 score ratings annotated by embryologists, which can be further divided into three grades: good, medium, and bad. | https://dataverse.harvard.edu/dataset.xhtml?persistentId=doi:10.7910/DVN/1Z4HLC |
| Adaptive adversarial neural networks for the analysis of lossy and domain-shifted datasets of medical images.[103] | 2022 | Static image | The dataset consists of 2440 images, including day 5 images of embryo development. The annotation is the five ratings 1-5, where ratings 1 and 2 are for no blastocysts developed and ratings 3,4 and 5 are for blastocysts developed. | https://osf.io/3kc2d/ |
| Embryo classification based on microscopic images.[104] | 2023 | Static image | The dataset consists of 1020 images, including images on day 3 and day 5 of embryo development. The annotations are good and bad ratings. | https://www.kaggle.com/competitions/world-championship-2023-embryo-classification/overview |
| Merging synthetic and real embryo data for advanced AI predictions.[105] | 2025 | Static image | The dataset consists of 5500 images generated by AI, including images from day 1-5 of embryo development. The annotations are the five grades 2-cell, 4-cell, 8-cell, morula, and blastocyst. | https://zenodo.org/records/14253170, https://huggingface.co/datasets/deepsynthbody/synembryo_latentdiffusion, https://huggingface.co/datasets/deepsynthbody/synembryo_stylegan |
| Automatic Identification of Human Blastocyst Components via Texture.[56] | 2017 | Static image | The dataset consists of 211 images, including day 5 images of embryo development. Annotations are segmented labels for different regions of the blastocyst, namely ICM, TE, BC, and ZP. | https://vault.sfu.ca/index.php/s/066vGJfviJMYuP6/authenticate |
| Cleavage-stage embryo segmentation using SAM-based dual branch pipeline: development and evaluation with the CleavageEmbryo dataset.[69] | 2024 | Static image | The dataset consists of 1548 images, including images from day 1 to 3 of embryo development. Annotation is the COCO object detection format, which detects objects separately for embryos, fragments, and background. | https://www.kaggle.com/datasets/austin012/cleavageembryo-dataset/data |
| Neural networks pipeline for quality management in IVF laboratory.[106] | 2024 | Table | The dataset consists of 11 1500 treatment cycles and includes 19 key tabular data. Label is the result of fetal heart clinical pregnancy. | https://github.com/embryossa/KAN-in-IVF |
| An annotated human blastocyst dataset to benchmark deep learning architectures for in vitro fertilization.[107] | 2023 | Static image and Table | The dataset consists of 2344 images from 837 cases, including images on day 4 or 5 of embryo development. These 2344 images were annotated with Expansion 5-grading, ICM 3-grading, and TE 3-grading. In addition, the 837 cases included six tabular features for hCG biochemical pregnancy, fetal heart clinical pregnancy, and live birth annotation. | https://figshare.com/articles/figure/Blastocyst_dataset_zip/20123153/3?file=39348899 |
IV-C Explainability
In ART, AI models are widely applied in embryo quality assessment, pregnancy success rate prediction and other key diagnosis and treatment tasks. Although these models show significant advantages in performance, their ”black box” feature makes explainability analysis be a prerequisite for clinical deployment. On the one hand, ART involves highly sensitive issues such as bioethics, patients’ right to informed consent, and the attribution of medical responsibility. The decision-making basis of the model must be transparent to doctors and patients. On the other hand, explanatory analysis not only helps enhance doctors’ trust in the prediction results of AI, but also helps identify potential medical-related features.
From the perspective of specific methods, t-distributed Stochastic Neighbor Embedding (t-SNE) [108] can be used for the visualization of high-dimensional feature spaces. By reducing the dimension of the embryo image features extracted by the deep model and mapping them to the two-dimensional plane, the clustering structure of embryo samples in the representation space can be revealed, which indirectly verifies whether the model has learned discriminative semantic features [37][38].
SHapley Additive exPlanations (SHAP) [109] can quantitatively analyze the predictive contributions of input variables, especially suitable for pregnancy prediction models constructed based on clinical tabular data, such as evaluating the positive and negative effects and weights of factors such as age, hormone levels, and embryo scores on the output probability [78][89].
Furthermore, for image-based model explainability, Gradient-weighted Class Activation Mapping (Grad-CAM) highlights the regions that the model is concerned about in the image by generating heat maps. Grad-CAM helps researchers to determine whether the AI has focused on clinically significant structures (such as ICM or TE) in embryonic images, thereby evaluating the rationality and usability of its decision-making process [35][110]. Overall, these interpretable technologies provide crucial support for AI applications in the ART field. They not only enhance the transparency of predictive models but also offer important clues for model optimization and clinical integration.
IV-D Public dataset
Since medical image data usually involve patient privacy information and legal compliance requirements, public datasets may violate relevant regulations or cannot pass ethical review, so most ART-related datasets have not been opened to the public. Considering the importance of open data to promote the development of medical artificial intelligence and computer-assisted therapy technology, this paper summarizes the currently publicly available representative datasets as listed in TABLE IV.
In addition to these public datasets, some open models such as STORK are also widely used for embryo evaluation. STORK (https://github.com/ih-lab/STORK) is an AI model from Stanford University, trained on the Inception image recognition architecture. The model can identify the morphological characteristics of the embryo and output the transfer probability or classification results. Other commercial software scoring systems, such as KIDScore and iDAScore [111], are also widely used. They are fully AI-driven scoring systems developed by Vitrolife, which are trained on time-lapse imaging sequences without manual labeling, outputting a continuous score (e.g., 1 to 10) for prioritization of embryos. Meanwhile, ERICA (Embryo Ranking Intelligent Classification Algorithm) [112] is also an embryo evaluation model based on deep learning, aiming to automatically score and rank static embryos in ART, which can be accessed on https://embryoranking.com/.
V Challenge and future work
The application of AI in ART has made remarkable progress in recent years, covering many key tasks from gamete quality assessment, embryo grading to pregnancy outcome prediction. Although previous studies have fully demonstrated the great potential of AI in improving assessment consistency, prediction accuracy and process automation, its clinical deployment still faces multiple challenges. In the following, the challenges of AI in the field of ART are systematically analyzed from four dimensions: multi-modal fusion, data resources, model generalization, and ethical barriers as shown in Fig. 4.
V-A Multi-modal fusion
In the process of ART, data types show a high degree of heterogeneity, including static image, time-lapse video, structured clinical data, and genomic information. Multi-modal data fusion is considered to be a key strategy to improve model prediction performance and achieve comprehensive embryo quality assessment. However, there are still several bottlenecks in the actual modeling process. First, there are differences in dimensionality and semantics between data, which makes it difficult to achieve effective fusion. So, the redundancy and even conflict in multi-modal data may reduce the accuracy of the model. Second, the complex structure of the fusion model leads to large computing resource consumption in the training process, and it is difficult to deploy in resource-constrained scenarios.
For the multi-modal fusion task, future research can focus on the design of lightweight fusion methods. For fusion strategies, in order to be able to effectively fuse the features among multi-modalities, common operation-based fusion methods such as AddFusion, ConcatFusion, and TensorFusion [113] can be used [114]. Furthermore, GNN, with their message passing mechanism based on graph structure, can establish effective connections between different modal features in multi-modal data, thereby mining high-order semantic relationships among cross-modal nodes and achieving information fusion [115]. At the same time, the Transformer architecture can effectively model the long-term dependence within a single modality through the self-attention mechanism, and its cross-attention mechanism further promotes the information interaction and alignment between different modalities [116]. The model can maintain the semantic consistency and complementarity between modalities in the fusion process. Compared with other fusion methods, decoupled fusion method decomposes multi-modal features into shared features and unique features, avoids redundant interference between modalities, and improves the model’s understanding of each modality [78][117][118][119].
The problem of lightweight model can be solved by using lightweight framework. For the image modality, MobileNet [120] and ShuffleNet [121] series networks can be used, and for the tabular modality, a simple MLP can be used. It can also be solved by knowledge distillation [122], using a large model (Teacher) to guide a small model (Student) to learn an intermediate representation or output distribution.
V-B Data resources
The data related to assisted reproduction have the characteristics of high sensitivity to human privacy, strict ethical approval requirements, and difficulty in sample collection, resulting in the general small datasets. And most of the experimental results reported in this paper are based on private datasets, which makes the results incomparable and hinders the technical development of AI in this field. In particular, there are still significant deficiencies in the construction of multi-modal datasets involving embryo morphological images and supporting clinical records. This problem directly limits the use of AI models in real-world scenarios.
In order to solve the problem of scarcity of public high-quality datasets in the field of ART, the most direct way is to collect clinical data manually. However, this method faces many institutional challenges in practice and is difficult to be widely promoted. In view of the difficulty of model training caused by the limited amount of data, a variety of effective alternative paths have been proposed in current research. Generative models such as GAN [123], Diffusion [124] can be used to synthesize datasets, thus increasing the diversity of datasets [105]. In addition, Few-shot Learning also provides an idea to deal with the small sample size problem. By introducing metric learning or meta-learning mechanisms, few-shot learning enables the model to achieve better performance with only a small number of labeled samples [125].
V-C Model generalization
Due to the differences in data acquisition equipment, scoring standards, and patient population characteristics between different laboratories, AI models often face the problem of performance degradation in the cross-institution deployment process, which exposes the limitations of insufficient generalization ability.
In order to improve the cross-domain adaptation ability of the model, the future should focus on the development of transfer learning [126] and domain adaptation [127] technology to realize the dynamic adjustment of model parameters or feature space. At the same time, as a collaborative training framework with local data, federated learning [128] provides a feasible path for breaking data islands and realizing cross-institution model sharing.
V-D Ethics and law
AI’s decision-making participation in ART, especially when it involves sensitive tasks such as embryo scoring, genetic screening, and pregnancy success prediction, raises a series of ethical and legal issues. For example, the scoring of embryos may be seen as judging the ”potential value of life”. In addition, AI intervention in the field of reproduction may also have a subtle impact on society’s conception of fertility and family structure ethics.
Therefore, it is urgent to establish a systematic ethical evaluation framework and technical supervision mechanism to clarify the responsibility of AI in assisted reproduction. And relevant practitioners should develop norms covering data privacy protection, informed consent, model interpretability, and transparency of use.
VI Conclusion
This review provides an in-depth discussion of the potential and challenges of artificial intelligence in embryo grading and pregnancy prediction in ART. By analyzing AI models from different data sources, we find that this technique not only effectively reduces the error caused by human factors, but also reveals the key factors affecting embryonic development through the processing of large amounts of data. However, there are still obstacles to the application of AI in ART, such as the difficulty of multi-modal data fusion and the scarcity of high-quality data. Future research should aim to address these issues and further optimize AI algorithms to adapt to different clinical needs. This will help to improve the success rate of IVF-ET, reduce the psychological stress of patients, and bring the hope of new life to more families.
Acknowledgment
This work is supported in part by the Guangdong Provincial Natural Science Foundation (2023A1515011431), the Guangzhou Science and Technology Planning Project (202201010092), the National Natural Science Foundation of China (72074105).